OpenStudy (anonymous):

How do I find the integral of (x^2+x+sqrt(x+1))/(x+1) dx

OpenStudy (anonymous):

A substitution hint would be maybe enough.

myininaya (myininaya):

\[\int\limits_{}^{}\frac{x(x+1)+\sqrt{x+1}}{\sqrt{x+1}} dx \] is that right ?

myininaya (myininaya):

try the sub u=sqrt(x+1) u^2=x+1 2udu=dx

OpenStudy (anonymous):

\[\int\frac{x^2+x+\sqrt{x+1}}{x+1}~dx=\int x~dx+\int(x+1)^{-1/2}~dx\]

OpenStudy (anonymous):

Thanks both!